# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ VI.2 Prospects for future work
Last, in addition to these avenues for the development of more accurate and generic models, there is also significant scope to further accelerate waveform generation with the few framework. As we identified in Section V.1, $\sim 90\%$ of the computational cost of the waveform derives from trajectory integration and waveform summation; there is scope to accelerate these two components in future work. The few integrator has been constructed entirely of vectorised array operations in preparation for batched trajectory integration on either CPU or GPU (depending on the batch size). While ODE integration is inherently serial, by computing derivatives for many trajectories in parallel the time-per-trajectory can be reduced significantly (potentially by orders of magnitude with GPUs and batch sizes $\gtrsim 10^{2}$). There are also multiple potential avenues towards improving the efficiency of the mode summation. While few does perform online mode selection, this operation is both expensive (requiring a full set of amplitudes to be computed) and sub-optimal (as can be verified with a slow, greedy-type minimisation of waveform mismatch). Improving this procedure may yield reduced computational cost without compromising the results of inference. Additionally, as is highlighted by the compact nature of each waveform mode in Fig. 1, time-frequency methods may significantly improve the mode summation for I/EMRI waveforms by more than an order of magnitude. This also presents an opportunity to integrate the application of the LISA response — itself an expensive operation — with this summation operation, directly computing TDI variables and further reducing computational costs. Development of these techniques for implementation in few is currently ongoing chapman-bird2025tf . Bringing together the myriad of potential improvements to be made to the efficiency of few, per-waveform wall-times of $\sim 1\,\mathrm{ms}$ are a realistic possibility, the achievement of which would greatly impact the identification and analysis of IMRI and EMRI signals.
@@ Acknowledgements.
We thank Ryuichi Fujita and Adam Pound for interesting discussions. We also thank Barry Wardell for support in setting up the few data download server, and Patrick Bourg and Christopher Whittall for improving the few tutorials. We are also indebted to BHPC (in particular Ryuichi Fujita, Hiroyuki Nakano, Norichika Sago and Masaru Shibata) for providing us with independent analytical and numerical data for the forcing functions and mode amplitudes used in Section IV and Section E.2 for the validation, upon request. We also thank the participants of the 2025 few Hackathon, hosted by the University of Southampton and funded by A. Pound’s Royal Society University Research Fellowship, for supporting key code developments in few and the production of this paper. CEAC-B acknowledges support from UKSA grant UKRI971. LS would like to acknowledge the support of the European Space Agency through ESA’s postdoctoral Research Fellowship programme. ZN acknowledges support from the ERC Consolidator/UKRI Frontier Research Grant GWModels (selected by the ERC and funded by UKRI [grant number EP/Y008251/1]). OB acknowledges support from the French space agency CNES in the framework of LISA. SK acknowledges the computing resources accessed from NUS IT Research Computing group and the support of the NUS Research Scholarship. JM and AJKC acknowledges support from the NUS Faculty of Science, under the research grant 22-5478-A0001. HK acknowledges the support from the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. Part of the analysis was performed on the “Symmetry” HPC at Perimeter Institute, using NVIDIA H200 GPU devices. JET and AJKC acknowledge support from the NASA LISA Preparatory Science grant 20-LPS20-0005. SI and AJKC acknowledge support from the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 A-800149200-00 (FY2023). SAH acknowledges support from NSF Grant PHY-2110384 and PHY-2409644, as well as the use of MIT Kavli Institute resources at MIT’s engaging cluster, subMIT resources of the MIT Department of Physics, and the Open Science Grid. NW acknowledges support from a Royal Society - Research Ireland University Research Fellowship. This publication has emanated from research conducted with the financial support of Research Ireland under Grant numbers 16/RS-URF/3428, 17/RS-URF-RG/3490 and 22/RS-URF-R/3825. This research was also done using services provided by the OSG Consortium osg07 ; osg09 ; OSPool ; OSDF , which is supported by the National Science Foundation awards #2030508 and #2323298.
@@ Appendix A Mass conventions
Inserting Eq. 23 into Eq. 18 and dropping all higher-order terms, we arrive at a new set of equations of motion, given by Eqs. 2 and 3. These equations are consistent with Eq. 18 at leading adiabatic order, but introduce new higher-order post-adiabatic corrections. Note that, based on Eq. 18, if we compute an adiabatic trajectory using our symmetric mass ratio convention, we can then relate it to an adiabatic trajectory produced via Eq. 18 by rescaling time by $m_{1}/M\times\nu/\epsilon$ and the phase by $\nu/\epsilon$. Alternatively, one can simply replace $m_{1}$ and $m_{2}$ by $M$ and $\mu$ in the small mass ratio model to produce the same trajectory as few.
@@ Appendix B Data grids and coordinate parameterizations
In our model, the simulation domain is divided into two regions: an inner region (Region A) and an outer region (Region B), following the approach in Bardwell:2025 . Region A is compact in the $p$-dimension and designed to densely sample amplitude and flux data near the separatrix, where these quantities exhibit rapid variation. To manage computational cost, we taper the maximum eccentricity as we approach the separatrix, thereby avoiding highly eccentric orbits near plunge, which are exponentially more expensive to simulate (see Section B.4 and Fig. 21). This tapering is also astrophysically motivated: gravitational radiation tends to circularize orbits, so we expect eccentricities to decrease as the separatrix is approached. Region B spans a much broader range in $p$ and $e$, and is sampled according to the post-Newtonian (PN) scaling behaviour of fluxes and amplitudes. In the following subsections, we detail the coordinate systems and sampling strategies used to construct our flux and amplitude datasets.
@@ B.1 Fluxes
00 0 $\leq e\leq S_{\mathrm{ecc}}(a,p,e;e_{\mathrm{sep}},e_{\mathrm{max}}),$ (25)
@@
providing the eccentricity tapering mentioned at the beginning of Appendix B. Furthermore, we have introduced the constants
@@
For this work, we set $x_{I}=1=y$, and to handle retrograde orbits, we set $a<0$ rather than switching the sign of $x_{I}$. Then we have $S_{\mathrm{ecc}}(a,p,e)=\tilde{S}_{\mathrm{gen}}(a,p,e,x_{I}=1)$. We compute the fluxes on a three-dimensional rectilinear grid by uniformly sampling the coordinates $(u,w,z)$ over the unit cube (i.e., $[0,1]^{3}$). The grid consists of $N_{u}\times N_{w}\times N_{z}$ points, with $(N_{u},N_{w},N_{z})=(129,65,65)$.
@@ B.2 Amplitude grid
For the amplitudes, we use the same bounds for Region A as we did for the fluxes (which are given in Eq. 27. To construct the inner amplitude grid, we also use the same coordinates as the fluxes, $(u,w,y,z)$ in Eq. 30, but with the modified constants
@@
with $\hat{U}_{p}(p)=p^{-1/2}$, $\hat{U}_{p,\mathrm{max}}(a,e,x_{I})=\hat{U}_{p}[\delta p^{B}_{\mathrm{min}}+p_{\mathrm{sep}}(a,e,x_{I})]$, $\hat{U}_{p,\mathrm{min}}(a,e,x_{I})=\hat{U}_{p}[\delta p^{B}_{\mathrm{max}}+p_{\mathrm{sep}}(a,e,x_{I})]$. Note that $(\hat{W},\hat{Y},\hat{Z})$ are identical to the outer grid coordinates $(W,Y,Z)$ used for the fluxes. To reduce the memory footprint of our interpolant by $37.5\%$, we reduce the sampling density with respect to Region A, such that $(N_{\hat{U}},N_{\hat{W}},N_{\hat{Z}})=(17,17,33)$ points in $(\hat{U},\hat{W},\hat{Z})$ — the density with respect to $Z$ remains fixed for the purposes of vectorised linear interpolation.
@@
Furthermore, we set $\delta p^{B}_{\mathrm{min}}=\delta p^{A}_{\mathrm{max}}$ for the fluxes to create a small region of overlap between A and B, which we used for preliminary cross-validation of interpolated flux values in the shared domain. We also found that using ${U}$ for the fluxes, instead of $\hat{U}$, and truncating Region B at $\delta p^{B}_{\mathrm{min}}$ and $\delta p^{B}_{\mathrm{max}}$ in (35) provided coordinates that much more closely align with the PN behaviour of the fluxes at large $p$. This change resulted in a modest but noticeable improvement in flux interpolation in the outer region.
@@ B.4 Timing of flux calculations
Based on the grids and sampling rates described above, we compute the fluxes at $615,810$ points in the $(a,p,e)$ parameter space. We perform these calculations on the Open Science Grid (OSG), a free open-access high throughput computing infrastructure osg07 ; osg09 ; OSPool ; OSDF . In Fig. 21 we plot the computational cost of each flux calculation for constant values of $a$. The computational cost is most significantly impacted by the value of the eccentricity and grows exponentially with $e$. However, this growth is not smooth due to the nature of the flux calculations performed by pybhpt. As the eccentricity grows, pybhpt needs to increase the numerical resolution of various computations, which it does by increasing the sampling rate by factors of two. Thus, at certain regions in parameter space, the calculation effectively doubles in computational cost because the calculation is hard coded to run at double the numerical resolution to meet precision requirements. This leads to the sharp horizontal bands in Fig. 21. Furthermore, particularly for large prograde spins, we observe a rapid growth in computational cost as we approach the separatrix. All together, the flux grid took just under 605,000 CPU hours to calculate.
Figure 21: The number of CPU minutes required to calculate the fluxes at each grid point for the same slices of $a$ visualized in Fig. 20.
@@ Appendix C Setup for Monte-Carlo studies of waveform characteristics
At multiple points throughout this work, we examine the accuracy and performance of our waveform model by performing Monte-Carlo draws of source parameters. For consistency, we randomly sample these parameters from the same distributions in each case; these are summarized in Table 3. Given a set of these drawn parameters, $p_{0}$ is chosen according to the desired time-to-plunge of the inspiral in question (typically either two or four years). For parameter sets where no such point can be found in our domain of validity (e.g., highly eccentric systems with $\epsilon\lesssim 10^{-6}$), a new point is drawn to replace it. This rejection procedure yields the more complicated distribution that appears in the top panel of Fig. 12.
For waveform-level comparisons, angular parameters are distributed in the typical manner: azimuthal and polar angles are distributed uniformly on the surface of the unit sphere, and phase parameters are uniformly distributed in the range $[0,2\pi)$. As none of our Monte-Carlo analyses entail random sampling of $d_{\mathrm{L}}$, we set this parameter to a fiducial value of $1\,\mathrm{Gpc}$ in all cases.
@@
$a$ 00 $0.999$ Uniform
$e_{0}$ 00 $0.9$ Uniform
Table 3: Distributions used for the Monte-Carlo studies of waveform characteristics performed in this work. Given a set of these drawn parameters, $p_{0}$ is chosen according to the desired time-to-plunge of the inspiral in question; see text of Appendix C for discussion.
Notably, Table 3 only covers prograde inspirals. The retrograde half of the parameter space is significantly less challenging to model accurately at all stages of waveform generation. We therefore exclude this region from many of our Monte-Carlo studies and focus primarily on prograde inspirals. None of the conclusions presented in this work change significantly when considering the retrograde case (to the contrary, the median performance of our model over the parameter space improves with such an extension).
@@ Appendix D GW data analysis framework
The PSD $S_{n}(f)$ describes the average mean-square fluctuations of $\tilde{n}(f)$ as a function of $f$. The diagonal noise covariance matrix defined in Eq. 52 gives rise to the Whittle likelihood function that is ubiquitous in the field of GW data analysis. The log of this likelihood function is
@@
The mismatch $\mathcal{M}$ is equal to zero for $h_{1}\propto h_{2}$ and equal to one if the two waveforms $h_{1}$ and $h_{2}$ are orthogonal. We will use Eq. 57 extensively in Section IV.3 as a figure of merit for the faithfulness of our waveform model.
@@
Equation 58 is optimal in the sense that it is the mean SNR (averaging over all noise realisations) under the assumption that the template waveform model is correct. Statistical biases in measured parameters due to noise fluctuations scale inversely with $\rho$, whereas systematic biases due to modelling errors do not; considering the range of expected SNRs for typical EMRIs is therefore an important facet of our analysis of waveform systematics in Section IV.3.
@@
where $\pi(\bm{\theta})$ is the prior probability, and the normalising constant $\mathcal{Z}$ is usually denoted the evidence (and is independent of $\bm{\theta}$). The posterior distribution $p(\bm{\theta}\mid d)$ encodes all information regarding the values of $\bm{\theta}$ that are best-supported by the observational data, given our prior expectations; insight regarding the form of $p(\bm{\theta}\mid d)$ is attained via parameter estimation. In Section V.6.1, we perform MCMC analyses (as implemented by the eryn package Karnesis:2023ras ; michael_katz_2023_7705496 ; Foreman-Mackey:2013 , with default settings) to sample EMRI posterior distributions, from the results of which we compute parameter variances as summary statistics. We assume wide uniform priors on all sampled parameters (such that they do not truncate the posterior bulk). Initial MCMC walker positions are chosen to be near the injected EMRI parameters, as our goal is not to perform search, only to probe the form of the posterior bulk. To incorporate the response function of LISA in our analyses, we use the fastlisaresponse package katz2022assessing to generate second-generation $\{A,E,T\}$ TDI variables Prince:2002hp from our EMRI waveforms, assuming spacecraft orbits with constant and equal arm-lengths. Our model for $S_{n}(f)$ is the SciRDv1 sensitivity curve babak2021lisa for this TDI variable convention, including the effects of the Galactic confusion noise PhysRevD.104.043019 as a stationary noise process robson2019construction .
Despite the computational efficiency of our model on GPUs, sampling Eq. 59 still typically takes $\sim 12\,\mathrm{hr}$ per EMRI source. This is prohibitively expensive for more exhaustive studies of parametric biases due to waveform model differences, such as those performed in Section V.3, in which a signal with parameters $\bm{\theta}_{\rm tr}$ is injected with one model and analysed with another, returning (potentially biased) best-fit parameters $\bm{\theta}_{\rm bf}$. To facilitate such an analysis, we estimate these biases via the linear signal approximation given by Cutler and Vallisneri Cutler:2007mi ,
@@
evaluated in Eq. 60 at $\bm{\theta}_{\rm tr}$.
Parameters estimated from noisy observations will be modified by a probabilistic shift away from the true parameter values. In a similar vein to Eq. 60, at leading order this shift is approximately
@@
where $n_{\mathrm{nr}}$ is a specific noise realisation. Since $n_{\mathrm{nr}}$ has zero mean, $\Delta\bm{\theta}_{\rm noise}$ is unbiased (i.e., $\mathbb{E}[\Delta\bm{\theta}_{\rm noise}]=0$). For all of the parameter estimation analyses we perform in this work — both in the form of Cutler-Vallisneri bias estimates (Section V.3) and MCMC explorations of posterior distributions (Section V.6) — we assume a “zero-noise” realisation, such that $n_{\mathrm{nr}}$ is the zero-vector. We make this choice to eliminate the statistical parameter shifts due to noise, leaving only the systematic shifts due to modelling errors (the characterisation of which is the goal of these analyses). Our results should therefore be interpreted as the expectation over all noise realisations. A pertinent question is why the posteriors recovered in Section V.6 have non-zero widths, despite the apparent absence of noise. It is important to emphasise that the “zero-noise” realisation is a perfectly valid draw from the distribution of all possible noise realisations, and that we should therefore expect our parameter estimates to be uncertain for zero noise according to our noise model (given by the PSD), just as we would expect to occur for any other realisation.
The linear signal approximation is typically valid for EMRIs with detectable SNRs $\rho\gtrsim 20$ (which gives rise to the Gaussian posterior distributions obtained in Section V.6). We use the StableEMRIFisher (SEF) package kejriwal_2024_sef to obtain robust information matrices given a set of EMRI parameters. Once the best-fit parameter point is obtained, we quantify its bias with respect to $\bm{\theta}_{\rm tr}$ by its corresponding sigma contour level. It is given by the Mahalanobis distance $D_{\rm Maha}$ mahalanobis , i.e. the multivariate generalization of the z-score, and is defined as
@@
We scale by $1/D$ here compared to the original definition since the unscaled Mahalanobis distance increases with dimensionality, which can lead to deceptively large sigma-contour levels. Consider, e.g., $\Gamma$ to be a $D\times D$ identity matrix and $\Delta\bm{\theta}$ a length $D$ vector of $1$’s, representing a $1\,\sigma$-biased point from the peak of a standard normal in $D$-dimensions. The numerator of Eq. 63 is $D$ in this case, justifying the scaling by $1/D$ to recover $D_{\rm Maha}=1$. Finally, note that parameter space correlations also impact the Mahalanobis distance through $\Gamma$.

### v2

@@ VI.2 Prospects for future work
Last, in addition to these avenues for the development of more accurate and generic models, there is also significant scope to further accelerate waveform generation with the few framework. As we identified in Section˜V.1, $\sim 90\%$ of the computational cost of the waveform derives from trajectory integration and waveform summation; there is scope to accelerate these two components in future work. The few integrator has been constructed entirely of vectorised array operations in preparation for batched trajectory integration on either CPU or GPU (depending on the batch size). While ODE integration is inherently serial, by computing derivatives for many trajectories in parallel the time-per-trajectory can be reduced significantly (potentially by orders of magnitude with GPUs and batch sizes $\gtrsim 10^{2}$). There are also multiple potential avenues towards improving the efficiency of the mode summation. While few does perform online mode selection, this operation is both expensive (requiring a full set of amplitudes to be computed) and sub-optimal (as can be verified with a slow, greedy-type minimisation of waveform mismatch). Improving this procedure may yield reduced computational cost without compromising the results of inference. Additionally, as is highlighted by the compact nature of each waveform mode in Fig.˜1, time-frequency methods may significantly improve the mode summation for I/EMRI waveforms by more than an order of magnitude. This also presents an opportunity to integrate the application of the LISA response — itself an expensive operation — with this summation operation, directly computing TDI variables and further reducing computational costs. Development of these techniques for implementation in few is currently ongoing chapman-bird2025tf . Bringing together the myriad of potential improvements to be made to the efficiency of few, per-waveform wall-times of $\sim 1\,\mathrm{ms}$ are a realistic possibility, the achievement of which would greatly impact the identification and analysis of IMRI and EMRI signals.
@@ Acknowledgements.
We thank Ryuichi Fujita and Adam Pound for interesting discussions. We also thank Barry Wardell for support in setting up the few data download server, and Patrick Bourg and Christopher Whittall for improving the few tutorials. We are also indebted to BHPC (in particular Ryuichi Fujita, Hiroyuki Nakano, Norichika Sago and Masaru Shibata) for providing us with independent analytical and numerical data for the forcing functions and mode amplitudes used in Section˜IV and Section˜E.2 for the validation, upon request. We also thank the participants of the 2025 few Hackathon, hosted by the University of Southampton and funded by A. Pound’s Royal Society University Research Fellowship, for supporting key code developments in few and the production of this paper. CEAC-B acknowledges support from UKSA grant UKRI971. LS would like to acknowledge the support of the European Space Agency through ESA’s postdoctoral Research Fellowship programme. ZN acknowledges support from the ERC Consolidator/UKRI Frontier Research Grant GWModels (selected by the ERC and funded by UKRI [grant number EP/Y008251/1]). OB acknowledges financial support from the Grant UKRI972 awarded via the UK Space Agency and computational resources from the French space agency CNES in the framework of LISA. SK acknowledges the computing resources accessed from NUS IT Research Computing group and the support of the NUS Research Scholarship. JM and AJKC acknowledges support from the NUS Faculty of Science, under the research grant 22-5478-A0001. HK acknowledges the support from the Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. Part of the analysis was performed on the “Symmetry” HPC at Perimeter Institute, using NVIDIA H200 GPU devices. JET and AJKC acknowledge support from the NASA LISA Preparatory Science grant 20-LPS20-0005. SI and AJKC acknowledge support from the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 A-800149200-00 (FY2023). SAH acknowledges support from NSF Grant PHY-2110384 and PHY-2409644, as well as the use of MIT Kavli Institute resources at MIT’s engaging cluster, subMIT resources of the MIT Department of Physics, and the Open Science Grid. NW acknowledges support from a Royal Society - Research Ireland University Research Fellowship. This publication has emanated from research conducted with the financial support of Research Ireland under Grant numbers 16/RS-URF/3428, 17/RS-URF-RG/3490 and 22/RS-URF-R/3825. This research was also done using services provided by the OSG Consortium osg07 ; osg09 ; OSPool ; OSDF , which is supported by the National Science Foundation awards #2030508 and #2323298.
@@ Appendix A Mass conventions
Inserting Eq.˜24 into Eq.˜19 and dropping all higher-order terms, we arrive at a new set of equations of motion, given by Eqs.˜2 and 3. These equations are consistent with Eq.˜19 at leading adiabatic order, but introduce new higher-order post-adiabatic corrections. Note that, based on Eq.˜19, if we compute an adiabatic trajectory using our symmetric mass ratio convention, we can then relate it to an adiabatic trajectory produced via Eq.˜19 by rescaling time by $m_{1}/M\times\nu/\epsilon$ and the phase by $\nu/\epsilon$. Alternatively, one can simply replace $m_{1}$ and $m_{2}$ by $M$ and $\mu$ in the small mass ratio model to produce the same trajectory as few.
@@ Appendix B Data grids and coordinate parameterizations
In our model, the simulation domain is divided into two regions: an inner region (Region A) and an outer region (Region B), following the approach in Bardwell:2025 . Region A is compact in the $p$-dimension and designed to densely sample amplitude and flux data near the separatrix, where these quantities exhibit rapid variation. To manage computational cost, we taper the maximum eccentricity as we approach the separatrix, thereby avoiding highly eccentric orbits near plunge, which are exponentially more expensive to simulate (see Section˜B.4 and Fig.˜21). This tapering is also astrophysically motivated: gravitational radiation tends to circularize orbits, so we expect eccentricities to decrease as the separatrix is approached. Region B spans a much broader range in $p$ and $e$, and is sampled according to the post-Newtonian (PN) scaling behaviour of fluxes and amplitudes. In the following subsections, we detail the coordinate systems and sampling strategies used to construct our flux and amplitude datasets.
@@ B.1 Fluxes
$0$ $\leq e\leq S_{\mathrm{ecc}}(a,p,e;e_{\mathrm{sep}},e_{\mathrm{max}}),$ (26)
@@
providing the eccentricity tapering mentioned at the beginning of Appendix˜B. Furthermore, we have introduced the constants
@@
For this work, we set $x_{I}=1=y$, and to handle retrograde orbits, we set $a<0$ rather than switching the sign of $x_{I}$. Then we have $S_{\mathrm{ecc}}(a,p,e)=\tilde{S}_{\mathrm{gen}}(a,p,e,x_{I}=1)$. We compute the fluxes on a three-dimensional rectilinear grid by uniformly sampling the coordinates $(u,w,z)$ over the unit cube (i.e., $0,1]^{3}$). The grid consists of $N_{u}\times N_{w}\times N_{z}$ points, with $(N_{u},N_{w},N_{z})=(129,65,65)$.
@@ B.2 Amplitude grid
For the amplitudes, we use the same bounds for Region A as we did for the fluxes (which are given in Eq.˜28. To construct the inner amplitude grid, we also use the same coordinates as the fluxes, $(u,w,y,z)$ in Eq.˜31, but with the modified constants
@@
[…]
@@
Furthermore, we set $\delta p^{B}_{\mathrm{min}}=\delta p^{A}_{\mathrm{max}}$ for the fluxes to create a small region of overlap between A and B, which we used for preliminary cross-validation of interpolated flux values in the shared domain. We also found that using U{U} for the fluxes, instead of $\hat{U}$, and truncating Region B at $\delta p^{B}_{\mathrm{min}}$ and $\delta p^{B}_{\mathrm{max}}$ in (36) provided coordinates that much more closely align with the PN behaviour of the fluxes at large $p$. This change resulted in a modest but noticeable improvement in flux interpolation in the outer region.
@@ B.4 Timing of flux calculations
Based on the grids and sampling rates described above, we compute the fluxes at $615,810$ points in the $(a,p,e)$ parameter space. We perform these calculations on the Open Science Grid (OSG), a free open-access high throughput computing infrastructure osg07 ; osg09 ; OSPool ; OSDF . In Fig.˜21 we plot the computational cost of each flux calculation for constant values of $a$. The computational cost is most significantly impacted by the value of the eccentricity and grows exponentially with $e$. However, this growth is not smooth due to the nature of the flux calculations performed by pybhpt. As the eccentricity grows, pybhpt needs to increase the numerical resolution of various computations, which it does by increasing the sampling rate by factors of two. Thus, at certain regions in parameter space, the calculation effectively doubles in computational cost because the calculation is hard coded to run at double the numerical resolution to meet precision requirements. This leads to the sharp horizontal bands in Fig.˜21. Furthermore, particularly for large prograde spins, we observe a rapid growth in computational cost as we approach the separatrix. All together, the flux grid took just under 605,000 CPU hours to calculate.
Figure 21: The number of CPU minutes required to calculate the fluxes at each grid point for the same slices of $a$ visualized in Fig.˜20.
@@ Appendix C Setup for Monte-Carlo studies of waveform characteristics
At multiple points throughout this work, we examine the accuracy and performance of our waveform model by performing Monte-Carlo draws of source parameters. For consistency, we randomly sample these parameters from the same distributions in each case; these are summarized in Table˜3. Given a set of these drawn parameters, $p_{0}$ is chosen according to the desired time-to-plunge of the inspiral in question (typically either two or four years). For parameter sets where no such point can be found in our domain of validity (e.g., highly eccentric systems with $\epsilon\lesssim 10^{-6}$), a new point is drawn to replace it. This rejection procedure yields the more complicated distribution that appears in the top panel of Fig.˜12.
For waveform-level comparisons, angular parameters are distributed in the typical manner: azimuthal and polar angles are distributed uniformly on the surface of the unit sphere, and phase parameters are uniformly distributed in the range $0,2\pi)$. As none of our Monte-Carlo analyses entail random sampling of $d_{\mathrm{L}}$, we set this parameter to a fiducial value of $1\,\mathrm{Gpc}$ in all cases.
@@
$a$ - 0 $0.999$ Uniform
$e_{0}$ - 0 $0.9$ Uniform
Table 3: Distributions used for the Monte-Carlo studies of waveform characteristics performed in this work. Given a set of these drawn parameters, $p_{0}$ is chosen according to the desired time-to-plunge of the inspiral in question; see text of Appendix˜C for discussion.
Notably, Table˜3 only covers prograde inspirals. The retrograde half of the parameter space is significantly less challenging to model accurately at all stages of waveform generation. We therefore exclude this region from many of our Monte-Carlo studies and focus primarily on prograde inspirals. None of the conclusions presented in this work change significantly when considering the retrograde case (to the contrary, the median performance of our model over the parameter space improves with such an extension).
@@ Appendix D GW data analysis framework
The PSD $S_{n}(f)$ describes the average mean-square fluctuations of $\tilde{n}(f)$ as a function of $f$. The diagonal noise covariance matrix defined in Eq.˜53 gives rise to the Whittle likelihood function that is ubiquitous in the field of GW data analysis. The log of this likelihood function is
@@
The mismatch $\mathcal{M}$ is equal to zero for $h_{1}\propto h_{2}$ and equal to one if the two waveforms $h_{1}$ and $h_{2}$ are orthogonal. We will use Eq.˜58 extensively in Section˜IV.3 as a figure of merit for the faithfulness of our waveform model.
@@
Equation˜59 is optimal in the sense that it is the mean SNR (averaging over all noise realisations) under the assumption that the template waveform model is correct. Statistical biases in measured parameters due to noise fluctuations scale inversely with $\rho$, whereas systematic biases due to modelling errors do not; considering the range of expected SNRs for typical EMRIs is therefore an important facet of our analysis of waveform systematics in Section˜IV.3.
@@
where $\pi(\bm{\theta})$ is the prior probability, and the normalising constant $\mathcal{Z}$ is usually denoted the evidence (and is independent of $\bm{\theta}$). The posterior distribution $p(\bm{\theta}\mid d)$ encodes all information regarding the values of $\bm{\theta}$ that are best-supported by the observational data, given our prior expectations; insight regarding the form of $p(\bm{\theta}\mid d)$ is attained via parameter estimation. In Section˜V.6.1, we perform MCMC analyses (as implemented by the eryn package Karnesis:2023ras ; michael_katz_2023_7705496 ; Foreman-Mackey:2013 , with default settings) to sample EMRI posterior distributions, from the results of which we compute parameter variances as summary statistics. We assume wide uniform priors on all sampled parameters (such that they do not truncate the posterior bulk). Initial MCMC walker positions are chosen to be near the injected EMRI parameters, as our goal is not to perform search, only to probe the form of the posterior bulk. To incorporate the response function of LISA in our analyses, we use the fastlisaresponse package katz2022assessing to generate second-generation $\{A,E,T\}$ TDI variables Prince:2002hp from our EMRI waveforms, assuming spacecraft orbits with constant and equal arm-lengths. Our model for $S_{n}(f)$ is the SciRDv1 sensitivity curve babak2021lisa for this TDI variable convention, including the effects of the Galactic confusion noise PhysRevD.104.043019 as a stationary noise process robson2019construction .
Despite the computational efficiency of our model on GPUs, sampling Eq.˜60 still typically takes $\sim 12\,\mathrm{hr}$ per EMRI source. This is prohibitively expensive for more exhaustive studies of parametric biases due to waveform model differences, such as those performed in Section˜V.3, in which a signal with parameters $\bm{\theta}_{\rm tr}$ is injected with one model and analysed with another, returning (potentially biased) best-fit parameters $\bm{\theta}_{\rm bf}$. To facilitate such an analysis, we estimate these biases via the linear signal approximation given by Cutler and Vallisneri Cutler:2007mi ,
@@
evaluated in Eq.˜61 at $\bm{\theta}_{\rm tr}$.
Parameters estimated from noisy observations will be modified by a probabilistic shift away from the true parameter values. In a similar vein to Eq.˜61, at leading order this shift is approximately
@@
where $n_{\mathrm{nr}}$ is a specific noise realisation. Since $n_{\mathrm{nr}}$ has zero mean, $\Delta\bm{\theta}_{\rm noise}$ is unbiased (i.e., $\mathbb{E}[\Delta\bm{\theta}_{\rm noise}]=0$). For all of the parameter estimation analyses we perform in this work — both in the form of Cutler-Vallisneri bias estimates (Section˜V.3) and MCMC explorations of posterior distributions (Section˜V.6) — we assume a “zero-noise” realisation, such that $n_{\mathrm{nr}}$ is the zero-vector. We make this choice to eliminate the statistical parameter shifts due to noise, leaving only the systematic shifts due to modelling errors (the characterisation of which is the goal of these analyses). Our results should therefore be interpreted as the expectation over all noise realisations. A pertinent question is why the posteriors recovered in Section˜V.6 have non-zero widths, despite the apparent absence of noise. It is important to emphasise that the “zero-noise” realisation is a perfectly valid draw from the distribution of all possible noise realisations, and that we should therefore expect our parameter estimates to be uncertain for zero noise according to our noise model (given by the PSD), just as we would expect to occur for any other realisation.
The linear signal approximation is typically valid for EMRIs with detectable SNRs $\rho\gtrsim 20$ (which gives rise to the Gaussian posterior distributions obtained in Section˜V.6). We use the StableEMRIFisher (SEF) package kejriwal_2024_sef to obtain robust information matrices given a set of EMRI parameters. Once the best-fit parameter point is obtained, we quantify its bias with respect to $\bm{\theta}_{\rm tr}$ by its corresponding sigma contour level. It is given by the Mahalanobis distance $D_{\rm Maha}$ mahalanobis , i.e. the multivariate generalization of the z-score, and is defined as
@@
We scale by $1/D$ here compared to the original definition since the unscaled Mahalanobis distance increases with dimensionality, which can lead to deceptively large sigma-contour levels. Consider, e.g., $\Gamma$ to be a $D\times D$ identity matrix and $\Delta\bm{\theta}$ a length $D$ vector of $1$’s, representing a $1\,\sigma$-biased point from the peak of a standard normal in $D$-dimensions. The numerator of Eq.˜64 is $D$ in this case, justifying the scaling by $1/D$ to recover $D_{\rm Maha}=1$. Finally, note that parameter space correlations also impact the Mahalanobis distance through $\Gamma$.